A cricket Ball of mass 150grms is moving with a velocity of 12m/s is hit by a bat so that the ball is turned back with a velocity of 20m/s the force of the blow acts for 0.01sec on the ball. The average force exerted by the bat on the ball is

Grade:10

3 Answers

Arun
6 years ago

Mass of the ball, m = 150 g = 0.15 kg

Initial speed of the ball, u = -12 m/s

Final speed of the bal after being hit, v = 20 m/s

Impulse = change in momentum = final momentum initial momentum

= Impulse = 0.15 × 20 0.15 × (-12) = 4.8 Ns

DHANAKONDA NARASIMHARAO
6 years ago
Why r u take initial velocity -ve. Impulse (I)=F×T. I=Ma×t =M(v-u) =0.15(20-12)=0.15×8=1.2NsI=1.2Ns
Arun
6 years ago
Initial velocity is in opposite direction.Hence I have taken this negative.RegardsArun(askIITians forum expert)